# Probing entanglement dynamics and topological transitions on noisy intermediate-scale quantum computers

### v2

@@ III.3 The twist operator
Figure 10: Results, obtained from ibm_auckland, for the twist order parameter for a chain with $L=8$ and PBC after quenches from the initial states $\left|{\psi_{\text{N}}}\right\rangle$ (a) and $\left|{\psi_{\text{S}}}\right\rangle$ (b).
Figure 11: The dynamical Berry phase for a chain with $L=8$ and PBC after a quench from the initial states $\left|{\psi_{\text{N}}}\right\rangle$, simulated on ibm_auckland.
@@ IV Summary and discussion
We have simulated quantum quenches in the SSH chain on NISQ devices provided by IBM. We considered particularly simple quench protocols in which the time evolution is governed by the topological SSH Hamiltonian in the fully dimerized limit. This time evolution can be implemented without the use of the Trotter decomposition. Our simulations on the IBM-Q devices demonstrated periodic oscillations in entanglement dynamics after a quench from a topologically trivial state, reflecting periodic transitions between two phases with different topological structures. We have also detected such dynamical topology using the twist order parameter and the Berry phase, which can be measured on a quantum device without the need for additional gates. Through some error mitigation techniques, we were able to deduce time-dependent data that match exact solutions for up to 12 qubits.
@@
The LSM twist operator applied in our simulations can formulate various quantities and concepts fundamental to quantum many-body physics, such as order parameters to characterize gapped phases [27, 46], Berry phases [36], polarization [36] and localization length [37]. Recently, the applications of the LSM twist operator have also been generalized to open quantum systems [47, 48]. The twist operator is diagonal in the computational basis on quantum computers. This property allows us to calculate its related quantities on classical computers by processing the bit-strings and their corresponding probabilities from quantum device measurements, without incurring additional overhead in quantum hardware. The method makes many interesting applications of the twist operator amenable to simulations on NISQ devices.
@@ Acknowledgements.
The authors acknowledge support from the IBM Q Hub at National Taiwan University, the National Science and Technology Counsil (NSTC) and the National Center for Theoretical Sciences (NCTS) in Taiwan. This work was supported under Grants No. NSTC 111-2119-M-007-009 and 112-2119-M-007-008.

### v3

@@ III.3 The twist operator
Figure 10: Results, obtained from ibm_auckland, for the twist order parameter for a chain with $L=8$ and PBC after quenches from the initial states $\left|{\psi_{\text{N}}}\right\rangle$ (a) and $\left|{\psi_{\text{S}}}\right\rangle$ (b). The dashed lines are results obtained by exact diagonalization (ED).
Figure 11: The dynamical Berry phase for a chain with $L=8$ and PBC after a quench from the initial state $\left|{\psi_{\text{N}}}\right\rangle$, simulated on ibm_auckland. The dashed lines are results obtained by exact diagonalization.
@@ IV Summary and discussion
We have simulated quantum quenches in the SSH chain on NISQ devices provided by IBM. We considered particularly simple quench protocols in which the time evolution is governed by the topological SSH Hamiltonian in the fully dimerized limit. This time evolution can be implemented without any Trotter error. Our simulations on the IBM-Q devices demonstrated periodic oscillations in entanglement dynamics after a quench from a topologically trivial state, reflecting periodic transitions between two phases with distinct topological structures. We have also detected such dynamical topology using the twist order parameter and the Berry phase, which can be measured on a quantum device without the need for additional gates. Through some error mitigation techniques, we were able to deduce time-dependent data that match exact solutions for up to 12 qubits.
@@
The LSM twist operator applied in our simulations can be used to formulate various quantities and concepts that are of fundamental importance in condensed matter physics, such as order parameters to characterize gapped phases [31, 54], Berry phases [42], polarization [42] and localization length [45]. Recently, the applications of the LSM twist operator have also been generalized to open quantum systems [55, 56]. The twist operator is diagonal in the computational basis on quantum computers. This property allows us to calculate its related quantities on classical computers by processing the bitstrings and their corresponding probabilities from quantum device measurements, without incurring additional overhead in quantum hardware. The method makes many interesting applications of the twist operator amenable to simulations on NISQ devices.
A natural extension of this work is to simulate the quench dynamics governed by the topological Hamiltonian with partial dimerization, i.e. with $(J^{\prime}>J\neq 0)$. In the partially dimerized case, the time evolution of the entanglement entropy has a more complex pattern. It exhibits slow oscillations with large amplitudes, which modulate faster and smaller oscillations [57]. One of the main challenges in simulating such quench dynamics using conventional Trotter decomposition on digital quantum computers is the need for deep quantum circuits to implement Trotterization, which often significantly reduces the accuracy of quantum simulations on current NISQ hardware. Nevertheless, various strategies have been proposed to reduce circuit depths for time evolution [11, 18, 58, 59, 60, 61]. It will be interesting to implement and extend these approaches on quantum computers to realize the complex dynamics of quantum entanglement and topology in the partially dimerized SSH model.
@@ Acknowledgements.
The authors acknowledge support from the IBM Q Hub at National Taiwan University, the National Science and Technology Council (NSTC) and the National Center for Theoretical Sciences (NCTS) in Taiwan. This work was supported under Grants No. NSTC 111-2119-M-007-009, 112-2119-M-007-008 and 112-2112-M-004-006.
Appendix A Derivation of the entanglement entropy
In this appendix, we present the analytical derivation of the entanglement entropies of the post-quench states.
A.1 Correlation matrix method for free-particle models
We first recall the correlation matrix method for calculating the entanglement entropy of a noninteracting system proposed by Peschel [38]. For a free-fermion system described by a quadratic Hamiltonian in terms of fermionic operators $c_{j}$ and $c_{j}^{\dagger}$, the reduced density matrix $\rho_{I}$ for a subsystem $I$ can be written in the form of a thermal density matrix with an entanglement Hamiltonian $\mathcal{H}_{I}$:
$\rho_{I}=\frac{1}{Z}e^{-\mathcal{H}_{I}},\quad\text{with }Z=\text{Tr}[e^{-\mathcal{H}_{I}}]\,,$ (27)
where the entanglement Hamiltonian is quadratic in the fermionic operators and can be diagonalized in the form:
$\mathcal{H}_{I}=\sum_{\ell\in I}\varepsilon_{\ell}f_{\ell}^{\dagger}f_{\ell}\,,$ (28)
with $c_{j}=\sum_{\ell}U_{j\ell}f_{\ell}$. The single-particle correlation function $C_{ij}=\langle{c_{i}^{\dagger}c_{j}}\rangle$ in subsystem $I$ can be constructed from the reduced density matrix $\rho_{I}$:
$C_{ij}=\text{Tr}[\rho_{I}c_{i}^{\dagger}c_{j}]=\sum_{\ell}U_{i\ell}^{*}U_{j\ell}\frac{1}{1+e^{\varepsilon_{\ell}}}\,.$ (29)
Thus, the correlation matrix $C$ has eigenvalues $\xi_{\ell}$ that are related to the eigenvalues $\varepsilon_{\ell}$ of the entanglement Hamiltonian via
$\xi_{\ell}=\frac{1}{1+e^{\varepsilon_{\ell}}}\,.$ (30)
The $\alpha$th-order Rényi entanglement entropy of the subsystem $I$, defined as
$S^{(\alpha)}=\frac{1}{1-\alpha}\log_{2}\text{Tr}[\rho_{I}^{\alpha}]\,,$ (31)
is explicitly given by
$S^{(\alpha)}=\frac{1}{1-\alpha}\sum_{\ell\in I}\left[\log_{2}(1+e^{-\alpha\varepsilon_{\ell}})-\log_{2}(1+e^{-\varepsilon_{\ell}})^{\alpha}\right]\,,$ (32)
or can be written in terms of the eigenvalues $\xi_{\ell}$ of the correlation matrix as
$S^{(\alpha)}=\frac{1}{1-\alpha}\sum_{\ell\in I}\log_{2}[(1-\xi_{\ell})^{\alpha}+\xi_{\ell}^{\alpha}]\,.$ (33)
For the case with $\alpha=2$, we obtain Eq. (11).
A.2 Time-dependent Wannier states
[…]
$|\psi_{k}(t)\rangle$ $=$ $\Big{\{}\left[\alpha\cos(d_{k}t)-i\beta\sin(d_{k}t)\hat{d}_{-}\right]a_{k}^{\dagger}$ (34)
$+$ $\left[\beta\cos(d_{k}t)-i\alpha\sin(d_{k}t)\hat{d_{+}}\right]b_{k}^{\dagger}\Big{\}}|\emptyset\rangle,$
where $d_{k}=|\mathbf{d}_{k}|$ and $\hat{d}_{\pm}$ is defined as $\hat{d}_{\pm}=(d_{k}^{x}\pm id_{k}^{y})/d_{k}=e^{\pm i\phi}$ with $\phi=\tan^{-1}(d_{k}^{y}/d_{k}^{x})$. In the flat band limit with $(J=0,J^{\prime}=1)$, we have $\hat{d}_{\pm}=e^{\pm ik}$; for the case with $(J=1,J^{\prime}=0)$, it is $\hat{d}_{\pm}=1$. Here we focus on the case $(J=0,J^{\prime}=1)$ as considered in this paper.
Applying the inverse Fourier transform to the state in Eq. (34), we obtain the Wannier state at the $m$th cell [62, 63],
$|w(m,t)\rangle$ $=$ $\frac{1}{\sqrt{N}}\sum_{k}e^{ikm}|\psi_{k}(t)\rangle$ (35)
$=$ $\Big{\{}\cos(d_{k}t)\left[\alpha a_{m}^{\dagger}+\beta b_{m}^{\dagger}\right]$
-- $i\sin(d_{k}t)\left[\beta a_{m_{+}}^{\dagger}+\alpha b_{m_{-}}^{\dagger}\right]\Big{\}}|\emptyset\rangle,$
where we have introduced $m_{\pm}=m\pm 1$ to denote the neighboring cells, and applied the relation $\frac{1}{N}\sum_{k}e^{ik(m-\ell)}=\delta_{\ell,m}$.
A.3 Quench from the Néel state $\left|{\psi_{\text{N}}}\right\rangle$
For the initial state $\left|{\psi_{\text{N}}}\right\rangle$ defined in Eq. (8), we have $\alpha=1,\beta=0$. The Wannier state after the quench is then
$|w_{\text{N}}(m,t)\rangle=\left[\cos(d_{k}t)a_{m}^{\dagger}-i\sin(d_{k}t)b_{m_{-}}^{\dagger}\right]|\emptyset\rangle\,.$ (36)
The correlation matrix associated with $\left|{w_{\text{N}}(m,t)}\right\rangle$ is given by
$\begin{pmatrix}\langle b_{m_{-}}^{\dagger}b_{m_{-}}\rangle&\langle b_{m_{-}}^{\dagger}a_{m}\rangle\\ \langle a_{m}^{\dagger}b_{m_{-}}\rangle&\langle a_{m}^{\dagger}a_{m}\rangle\end{pmatrix}=\begin{pmatrix}\sin^{2}(d_{k}t)&-i\sin(d_{k}t)\cos(d_{k}t)\\ i\sin(d_{k}t)\cos(d_{k}t)&\cos^{2}(d_{k}t)\end{pmatrix}.$ (37)
The eigenvalues of this correlation matrix are 00 and $1$, resulting in zero EE. This indicates that the Wannier state does not contribute to the EE if it lies entirely within the subsystem. Here we consider a symmetric bipartition with a boundary between the $(m-1)$-th and $m$-th cells. The submatrix of the left half is $\langle{b_{m_{-}}^{\dagger}b_{m_{-}}}\rangle$, with an eigenvalue $\sin^{2}(d_{k}t)$. The second-order Rényi EE of the left half is then
$S_{\text{OBC}}$ $=$ $-\log_{2}\left[\sin^{4}(d_{k}t)+\cos^{4}(d_{k}t)\right]$ (38)
$=$ $-\log_{2}\left[1-\sin^{2}(2d_{k}t)/2\right]\,,$
where the subscript ’OBC’ denotes open boundary conditions. In a chain with PBC, there are two boundaries between the partitions, with each boundary intersecting a Wannier state. Consequently, the Rényi EE for PBC is twice that of OBC.
$S_{\text{PBC}}=-2\log_{2}\left[1-\sin^{2}(2d_{k}t)/2\right]\,.$ (39)
A.4 Quench from the fully dimerized state $\left|{\psi_{\text{S}}}\right\rangle$
For a quench from the fully dimerized state $\left|{\psi_{\text{S}}}\right\rangle$, the post-quench Wannier state with $\alpha=1/\sqrt{2},\beta=-1/\sqrt{2}$ becomes
$|w_{\text{S}}(m,t)\rangle$ $=$ $\frac{1}{\sqrt{2}}\Big{\{}\cos(d_{k}t)\left[a_{m}^{\dagger}-b_{m}^{\dagger}\right]$ (40)
$+$ $i\sin(d_{k}t)\left[a_{m_{+}}^{\dagger}-b_{m_{-}}^{\dagger}\right]\Big{\}}|\emptyset\rangle\,,$
where the second term involves two next-nearest-neighbor unit cells.
We consider a bipartition with a boundary located between the unit cells $m$ and $m+1$. Two Wannier states are associated with this boundary: one formed between cells $m-1$ and $m$, and the other between cells $m$ and $m+2$. Thus, we need to include these two Wannier states $|w_{\text{S}}(m,t)\rangle$ and $|w_{\text{S}}(m+1,t)\rangle$ to construct the relevant portion of the correlation matrix:
$\begin{pmatrix}\langle b_{m_{-}}^{\dagger}b_{m_{-}}\rangle&\langle b_{m_{-}}^{\dagger}a_{m}\rangle&\langle b_{m_{-}}^{\dagger}b_{m}\rangle\\ \langle a_{m}^{\dagger}b_{m_{-}}\rangle&\langle a_{m}^{\dagger}a_{m}\rangle&\langle a_{m}^{\dagger}b_{m}\rangle\\ \langle b_{m}^{\dagger}b_{m_{-}}\rangle&\langle b_{m}^{\dagger}a_{m}\rangle&\langle b_{m}^{\dagger}b_{m}\rangle\end{pmatrix}=\frac{1}{2}\left(\begin{array}[]{cccc}{\sin^{2}(d_{k}/t)}&-i\sin(d_{k}t)\cos(d_{k}t)&i\sin(d_{k}t)\cos(d_{k}t)\\ i\sin(d_{k}t)\cos(d_{k}t)&{\cos^{2}(d_{k}t)}&-\cos^{2}(d_{k}t)\\ -i\sin(d_{k}t)\cos(d_{k}t)&-\cos^{2}(d_{k}t)&1\\ \end{array}\right).$ (44)
The eigenvalues of this matrix are $20,(1-\cos(d_{k}t))/$ 2 and $(1+\cos(d_{k}t))/2$. Thus, the second-order Rényi EE with a single partition boundary, as in an open chain, is given by
$S_{\text{OBC}}=-2\log_{2}\left[1-\sin^{2}(d_{k}t)/2\right].$ (45)
In a chain with PBC, two boundaries separate the partitions, doubling the Rényi EE:
$S_{\text{PBC}}=-4\log_{2}\left[1-\sin^{2}(d_{k}t)/2\right].$ (46)